# DELVE & MAGIC Collaborations Spectroscopic Analysis of Pictor II: a very low metallicity ultra-faint dwarf galaxy bound to the Large Magellanic Cloud⋆

### v1

@@ 4.3 Detailed abundances of an extremely metal-poor star
Carbon-enhanced metal-poor stars are thought to be especially related to Pop III stars (e.g., Frebel & Norris, 2015). About 40% of halo stars with have (Placco et al., 2014; Arentsen et al., 2022), but it is still debated whether a similar trend is found in dwarf galaxies like Sculptor (e.g., Salvadori et al., 2015; Chiti et al., 2018a) and Sagittarius (e.g., Limberg et al., 2023; Sestito et al., 2024). We measure this Pic II star to have , and corrections for internal mixing in red giant evolution (Placco et al., 2014) suggest a natal abundance . This star is thus clearly not significantly carbon-enhanced. Including this star, there are now 22 stars across 9 UFDs that have and carbon measurements: 3 in Segue 1, (Frebel et al., 2014); 2 in UMa II, (Frebel et al., 2010); 3 in Bootes I, (Norris et al., 2010; Gilmore et al., 2013; Frebel et al., 2016); 3 in Ret II; (Ji et al., 2016c; Roederer et al., 2016); 6 in Tuc II, (Ji et al., 2016d; Chiti et al., 2018b, 2023); 2 in Car II (Ji et al., 2020b); 1 in Car III (Ji et al., 2020b); and 1 in Tuc V (Hansen et al., 2024). 13 (8) of these 22 stars are carbon-enhanced with (), which is higher than but consistent with the stellar halo fraction, as shown in Ji et al. (2020b). We note that every one of these UFDs other than Pic II has at least one carbon-enhanced star at , but with one star in Pic II it would be premature to conclude that Pic II does not contain any carbon-enhanced stars (e.g., Koch et al., 2008; Gilmore et al., 2013; Ishigaki et al., 2014; Ji et al., 2016d, 2019, 2020b; Chiti et al., 2023; Waller et al., 2023; Webber et al., 2023; Hansen et al., 2024).
@@
The -elements (magnesium, silicon, calcium, titanium) in this Pic II star have super-solar ratios, similar to stars in all UFDs other than Hor I. Nagasawa et al. (2018) found three stars in that system with . As Hor I is very likely associated with the LMC (Koposov et al., 2015b), those authors hypothesized that the low -element abundances could be an indication of association with the LMC, possibly due to an overall lower star formation efficiency for galaxies forming near the LMC instead of the MW. However, both Car II (Ji et al., 2020b) and Ret II (Ji et al., 2016c; Roederer et al., 2016), two other LMC UFDs, do not show this trend. Pic II provides additional evidence against that hypothesis, as it is also clearly associated with the LMC but contains a star with enhanced [/Fe] ratios. However, this Pic II star has , significantly lower than the stars studied so far in Hor I. This low-[Fe/H] star does not probe the late-time chemical evolution of the system, and so cannot really test the hypothesis that star formation efficiencies are different for LMC satellites. Unfortunately, we currently have no -element constraints in Pic II stars of higher [Fe/H].
@@
Ratios of different -elements provide some constraint on the progenitor mass of a core-collapse supernova (e.g., McWilliam et al., 2013; Carlin et al., 2018; Ji et al., 2020b). In particular, the yield of hydrostatically synthesized elements like Mg increases significantly with progenitor mass, while explosively synthesized elements like Ca are less affected by progenitor mass. The detailed relation between [Mg/Ca] and progenitor mass is not straightforward, as Ca yields are affected by the SN explosion, but higher [Mg/Ca] ratios correspond to higher progenitor masses on average (e.g., Heger & Woosley, 2010). We measure in this star, a high ratio that suggests this Fe-poor star is predominantly (or perhaps only) enriched by relatively massive core-collapse supernovae.
@@
Ji et al. (2020b) explored the evolution of [Mg/Ca] versus [Fe/H] for the UFD population and while most UFDs do not show a trend with [Fe/H], two LMC satellites (Car II and Ret II), show a declining trend with increasing [Fe/H]. The lowest metallicity star in Car III also has high [Mg/Ca] but there are only two stars in Car III with chemical abundance measurements. This trend and its difference from other non-LMC UFDs could be evidence for environmental effects in the chemical evolution of a UFD (see Section 4.3.3 of Ji et al., 2020b). However, Hansen et al. (2024) recently showed that Tucana V, which does not appear to be associated with the LMC, also exhibits a similar [Mg/Ca] trend as the Magellanic satellites identified by Ji et al. (2020b). The star formation histories of LMC UFDs suggest they formed later (Sacchi et al., 2021; Simon et al., 2023; Durbin et al., 2025). While we cannot determine how [Mg/Ca] evolves with [Fe/H] in Pic II, the high value of [Mg/Ca]=0.51 at [Fe/H]=-3.3 suggests Pic II may follow the trend observed in other LMC UFDs.
@@
The main other element of note is scandium. We measured from 7 different lines, an unusually high value compared to other UFDs and halo stars at similar [Fe/H]. We fit synthetic spectra including hyperfine splitting for all Sc lines, the line-to-line abundance scatter is low, and variations due to stellar parameters are included in the uncertainty. We also verify that our analysis method reproduces the Sc abundance of several other UFD and standard halo stars to within 0.1 dex. The somewhat high Sc abundance in this star is thus real and significant, at least in 1D LTE (3D and NLTE effects for Sc II have not been investigated in metal-poor red giant stars, Bergemann & Nordlander 2014). The interpretation of high Sc is somewhat unclear, as standard models of core-collapse supernovae tend to underproduce Sc (see discussion in Nomoto et al., 2013; Curtis et al., 2019). Two ways of increasing Sc production include the -process (Fröhlich et al., 2006) or high-energy hypernova (Nomoto et al., 2006), but the other elements available in our Pic II star do not constrain these processes. We note that Zn tends to be produced with Sc in these models, but our spectrum provides only a weak 3 (5) limit of . Overall, it seems that there could be substantial [Sc/Fe] variation across different UFDs (e.g., Ji et al., 2020b; Chiti et al., 2023), which may point to interesting variations in the first supernova explosions.
@@ 4.4 Astrophysical J-factor and D-factor
The MW UFDs are among the most promising targets for searches of the products of dark matter annihilation or decay as they are nearby, highly dark matter-dominated, and nearly background free (e.g., Ackermann et al., 2015; Geringer-Sameth et al., 2015a; McDaniel et al., 2024; Circiello et al., 2025). While some individual UFDs are excellent targets on their own, new UFDs are also useful in stacked joint analyses to boost any signal or exclusion limits (e.g., Ackermann et al., 2011; Geringer-Sameth & Koushiappas, 2011) and we analyze Pic II here for future dark matter indirect detection studies.
@@ 5 Summary
The Legacy Survey of Space and Time (Ivezić et al., 2019) at the upcoming Vera C. Rubin Observatory will find many more UFDs and potentially more LMC associated systems (e.g., Tsiane et al., 2025). Identifying LMC associated UFDs will enable the study of UFDs that formed in a lower density environment. Stellar spectroscopic follow-up of UFD candidates is key to classify and confirm these systems and to determine their kinematics, orbital properties, metallicity distributions, and dark matter properties. The next generation of instruments (e.g., VLT/MOONS181818https://vltmoons.org/, VIA191919https://via-project.org/, Subaru Prime Focus Spectrograph202020https://pfs.ipmu.jp), ongoing and upcoming surveys (e.g., DESI, 4MOST), and the next generation of 30 m-class telescopes will be key to characterize the distant UFDs that will be found with Legacy Survey of Space and Time, Euclid, and the Nancy Grace Roman Space Telescope.
@@ Acknowledgments
We thank Andrew Casey and Anna Frebel for co-developing the latest version of SMH. We thank Chris Garling for useful discussions. APJ was supported by NASA through Hubble Fellowship grant HST-HF2-51393.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. W.C. gratefully acknowledges support from a Gruber Science Fellowship at Yale University. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2139841. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. AC is supported by a Brinson Prize Fellowship at UChicago/KICP.
@@
astropy (Astropy Collaboration et al., 2013, 2018), matplotlib (Hunter, 2007), NumPy (Walt et al., 2011), iPython (Pérez & Granger, 2007), SciPy (Virtanen et al., 2020) corner.py (Foreman-Mackey, 2016), emcee (Foreman-Mackey et al., 2013), gala (Price-Whelan, 2017), galpy212121 http://github.com/jobovy/galpy (Bovy, 2015), CarPy (Kelson, 2003), MOOG (Sneden, 1973; Sobeck et al., 2011) SMH (Casey, 2014) pandas (pandas development team, 2020), seaborn (Waskom et al., 2016), pocoMC (Karamanis et al., 2022a, b), NOIRLab IRAF (Tody, 1986, 1993; Fitzpatrick et al., 2024)
@@ Appendix B Literature references for figures
Our references are drawn from the Local Volume Database (Pace, 2024)222222Release 1.0.5 https://github.com/apace7/local_volume_database. References for UFDs including the LMC systems (Bellazzini et al., 2005; Belokurov et al., 2007; Bhardwaj et al., 2024; Boettcher et al., 2013; Bruce et al., 2023; Cantu et al., 2021; Carlin et al., 2009, 2017; Carlin & Sand, 2018; Casey et al., 2025; Cerny et al., 2021a, b, 2023a, 2023b, 2025; Chiti et al., 2021, 2022, 2023; Cicuéndez et al., 2018; Correnti et al., 2009; Crnojević et al., 2016; Dall’Ora et al., 2006, 2012; Drlica-Wagner et al., 2015; Fritz et al., 2019; Garofalo et al., 2013, 2025; Greco et al., 2008; Hansen et al., 2024; Heiger et al., 2024; Homma et al., 2018, 2019, 2024; Jenkins et al., 2021; Ji et al., 2021; Karczmarek et al., 2015; Kim & Jerjen, 2015b; Kim et al., 2016b; Kirby et al., 2013a, 2015, 2017; Koposov et al., 2011, 2015a, 2015b, 2018; Kuehn et al., 2008; Lee et al., 2009; Li et al., 2017, 2018a; Longeard et al., 2018; Martínez-Vázquez et al., 2015, 2019, 2021a, 2021b; Mateo et al., 2008; McConnachie, 2012; Medina et al., 2018; Moskowitz & Walker, 2020; Muñoz et al., 2018b; Musella et al., 2009; Mutlu-Pakdil et al., 2018, 2020; Oakes et al., 2022; Pace et al., 2020; Richstein et al., 2022, 2024; Sand et al., 2012; Simon & Geha, 2007; Simon et al., 2011, 2015, 2017; Simon, 2019; Simon et al., 2020; Smith et al., 2023; Spencer et al., 2017, 2018; Stetson et al., 2014; Tan et al., 2025; Torrealba et al., 2016a, b, 2018; Vivas et al., 2016, 2020, 2022; Walker et al., 2009, 2015b, 2015a; Walsh et al., 2008; Wang et al., 2019; Willman et al., 2006, 2011). References for ambiguous systems (Balbinot et al., 2013; Cerny et al., 2023b, c; Conn et al., 2018; Gatto et al., 2022; Homma et al., 2019; Kim & Jerjen, 2015a; Kim et al., 2015, 2016a; Luque et al., 2018; Martin et al., 2016; Mau et al., 2019, 2020; Muñoz et al., 2018b; Simon et al., 2024; Smith et al., 2024; Torrealba et al., 2019). References for the GC systems (Alonso-García et al., 2025; Bandyopadhyay et al., 2025; Baumgardt & Hilker, 2018; Baumgardt et al., 2020; Baumgardt & Vasiliev, 2021; Bradford et al., 2011; Çalışkan et al., 2012; Carballo-Bello et al., 2016; Crociati et al., 2023; Da Costa et al., 1992; Deras et al., 2023; Fanelli et al., 2024; Frelijj et al., 2025; Geisler et al., 2023; Gieles et al., 2021; González-Díaz et al., 2023; Hamren et al., 2013; Hankey & Cole, 2011; Harris, 1996; Henao et al., 2025; Johnson et al., 2018; Joo et al., 2019; Jordi et al., 2009; Kirby et al., 2015; Kobulnicky et al., 2005; Koch et al., 2009; Kunder & Butler, 2020; Kunder et al., 2021, 2024; Kurtev et al., 2008; Latour et al., 2025; Leanza et al., 2024; Longeard et al., 2021; Loriga et al., 2025; Montecinos et al., 2021; Muñoz et al., 2018a, b; Pallanca et al., 2023; Richstein et al., 2024; Rosenberg et al., 1998; Saviane et al., 2012; Simpson, 2018; Souza et al., 2021; Strader & Kobulnicky, 2008; Vásquez et al., 2018; Voggel et al., 2016; Weisz et al., 2016).
@@
The chemical abundance references of the UFD stars in Figure 9 are as follows: Boötes I (Gilmore et al., 2013; Frebel et al., 2016; Waller et al., 2023), Boötes II (Ji et al., 2016b), Canes Venatici II (François et al., 2016), Carina II and III (Ji et al., 2020a), Cetus II (Webber et al., 2023), Grus I (Ji et al., 2019), Grus II (Hansen et al., 2020), Coma Berenices and Ursa Major II (Frebel et al., 2010), Hercules (Koch et al., 2008; François et al., 2016), Horologium I (Nagasawa et al., 2018), Leo IV (Simon et al., 2010; François et al., 2016), Pisces II (Spite et al., 2018), Reticulum II (Ji et al., 2016a, c) Segue 1 (Frebel et al., 2014), Segue 2 (Roederer & Kirby, 2014), Triangulum II (Venn et al., 2017; Kirby et al., 2017; Ji et al., 2019), Tucana II (Chiti et al., 2018b, 2023), Tucana III (Hansen et al., 2017; Marshall et al., 2019), Tucana V (Hansen et al., 2024), and Ursa Major I (Waller et al., 2023).

### v2

@@ 4.3 Detailed abundances of an extremely metal-poor star
Carbon-enhanced metal-poor stars are thought to be especially related to Pop III stars (e.g., Frebel & Norris, 2015). About 40% of halo stars with have (Placco et al., 2014; Arentsen et al., 2022), but it is still debated whether a similar trend is found in dwarf galaxies like Sculptor (e.g., Salvadori et al., 2015; Chiti et al., 2018b) and Sagittarius (e.g., Limberg et al., 2023; Sestito et al., 2024). We measure this Pic II star to have , and corrections for internal mixing in red giant evolution (Placco et al., 2014) suggest a natal abundance . This star is thus clearly not significantly carbon-enhanced. Including this star, there are now 22 stars across 9 UFDs that have and carbon measurements: 3 in Segue 1, (Frebel et al., 2014); 2 in UMa II, (Frebel et al., 2010); 3 in Bootes I, (Norris et al., 2010; Gilmore et al., 2013; Frebel et al., 2016); 3 in Ret II; (Ji et al., 2016c; Roederer et al., 2016); 6 in Tuc II, (Ji et al., 2016b; Chiti et al., 2018a, 2023); 2 in Car II (Ji et al., 2020a); 1 in Car III (Ji et al., 2020a); and 1 in Tuc V (Hansen et al., 2024). 13 (8) of these 22 stars are carbon-enhanced with (), which is higher than but consistent with the stellar halo fraction, as shown in Ji et al. (2020a). We note that every one of these UFDs other than Pic II has at least one carbon-enhanced star at , but with one star in Pic II it would be premature to conclude that Pic II does not contain any carbon-enhanced stars (e.g., Koch et al., 2008; Gilmore et al., 2013; Ishigaki et al., 2014; Ji et al., 2016b, 2019, 2020a; Chiti et al., 2023; Waller et al., 2023; Webber et al., 2023; Hansen et al., 2024).
@@
The -elements (magnesium, silicon, calcium, titanium) in this Pic II star have super-solar ratios, similar to stars in all UFDs other than Hor I. Nagasawa et al. (2018) found three stars in that system with . As Hor I is very likely associated with the LMC (Koposov et al., 2015b), those authors hypothesized that the low -element abundances could be an indication of association with the LMC, possibly due to an overall lower star formation efficiency for galaxies forming near the LMC instead of the MW. However, both Car II (Ji et al., 2020a) and Ret II (Ji et al., 2016c; Roederer et al., 2016), two other LMC UFDs, do not show this trend. Pic II provides additional evidence against that hypothesis, as it is also clearly associated with the LMC but contains a star with enhanced [/Fe] ratios. However, this Pic II star has , significantly lower than the stars studied so far in Hor I. This low-[Fe/H] star does not probe the late-time chemical evolution of the system, and so cannot really test the hypothesis that star formation efficiencies are different for LMC satellites. Unfortunately, we currently have no -element constraints in Pic II stars of higher [Fe/H].
@@
Ratios of different -elements provide some constraint on the progenitor mass of a core-collapse supernova (e.g., McWilliam et al., 2013; Carlin et al., 2018; Ji et al., 2020a). In particular, the yield of hydrostatically synthesized elements like Mg increases significantly with progenitor mass, while explosively synthesized elements like Ca are less affected by progenitor mass. The detailed relation between [Mg/Ca] and progenitor mass is not straightforward, as Ca yields are affected by the SN explosion, but higher [Mg/Ca] ratios correspond to higher progenitor masses on average (e.g., Heger & Woosley, 2010). We measure in this star, a high ratio that suggests this Fe-poor star is predominantly (or perhaps only) enriched by relatively massive core-collapse supernovae. Non-LTE corrections corrections171717Online correction tools at https://spectrum.inasan.ru/nLTE/ for Ca, https://www.inspect-stars.com for Mg, and https://nlte.mpia.de for both. for this star result in small increases in Mg ( dex, Osorio et al. 2015; Bergemann et al. 2017) and moderate increases in Ca ( dex, Mashonkina et al. 2007, 2016), so the [Mg/Ca] ratio is 0.2 dex lower but still significantly enhanced.
@@
Ji et al. (2020a) explored the evolution of [Mg/Ca] versus [Fe/H] for the UFD population and while most UFDs do not show a trend with [Fe/H], two LMC satellites (Car II and Ret II), show a declining trend with increasing [Fe/H]. The lowest metallicity star in Car III also has high [Mg/Ca] but there are only two stars in Car III with chemical abundance measurements. This trend and its difference from other non-LMC UFDs could be evidence for environmental effects in the chemical evolution of a UFD (see Section 4.3.3 of Ji et al., 2020a). However, Hansen et al. (2024) recently showed that Tucana V, which does not appear to be associated with the LMC, also exhibits a similar [Mg/Ca] trend as the Magellanic satellites identified by Ji et al. (2020a). The star formation histories of LMC UFDs suggest they formed later (Sacchi et al., 2021; Simon et al., 2023; Durbin et al., 2025). While we cannot determine how [Mg/Ca] evolves with [Fe/H] in Pic II, the high value of at suggests Pic II may follow the trend observed in other LMC UFDs.
@@
The main other element of note is scandium. We measured from 7 different lines, an unusually high value compared to other UFDs and halo stars at similar [Fe/H]. We fit synthetic spectra including hyperfine splitting for all Sc lines, the line-to-line abundance scatter is low, and variations due to stellar parameters are included in the uncertainty. We also verify that our analysis method reproduces the Sc abundance of several other UFD and standard halo stars to within 0.1 dex. The somewhat high Sc abundance in this star is thus real and significant, at least in 1D LTE (3D and NLTE effects for Sc II have not been investigated in metal-poor red giant stars, Bergemann & Nordlander 2014). The interpretation of high Sc is somewhat unclear, as standard models of core-collapse supernovae tend to underproduce Sc (see discussion in Nomoto et al., 2013; Curtis et al., 2019). Two ways of increasing Sc production include the -process (Fröhlich et al., 2006) or high-energy hypernova (Nomoto et al., 2006), but the other elements available in our Pic II star do not constrain these processes. We note that Zn tends to be produced with Sc in these models, but our spectrum provides only a weak 3 (5) limit of . Overall, it seems that there could be substantial [Sc/Fe] variation across different UFDs (e.g., Ji et al., 2020a; Chiti et al., 2023), which may point to interesting variations in the first supernova explosions.
@@ 4.4 Astrophysical J-factor and D-factor
The MW UFDs are among the most promising targets for searches of the products of dark matter annihilation or decay as they are nearby, highly dark matter-dominated, and nearly background free (e.g., Ackermann et al., 2015; Geringer-Sameth et al., 2015b; McDaniel et al., 2024; Circiello et al., 2025). While some individual UFDs are excellent targets on their own, new UFDs are also useful in stacked joint analyses to boost any signal or exclusion limits (e.g., Ackermann et al., 2011; Geringer-Sameth & Koushiappas, 2011) and we analyze Pic II here for future dark matter indirect detection studies.
@@ 5 Summary
The Legacy Survey of Space and Time (Ivezić et al., 2019) at the upcoming Vera C. Rubin Observatory will find many more UFDs and potentially more LMC associated systems (e.g., Tsiane et al., 2025). Identifying LMC associated UFDs will enable the study of UFDs that formed in a lower density environment. Stellar spectroscopic follow-up of UFD candidates is key to classify and confirm these systems and to determine their kinematics, orbital properties, metallicity distributions, and dark matter properties. The next generation of instruments (e.g., VLT/MOONS191919https://vltmoons.org/, VIA202020https://via-project.org/, Subaru Prime Focus Spectrograph212121https://pfs.ipmu.jp), ongoing and upcoming surveys (e.g., DESI, 4MOST), and the next generation of 30 m-class telescopes will be key to characterize the distant UFDs that will be found with Legacy Survey of Space and Time, Euclid, and the Nancy Grace Roman Space Telescope.
@@ Acknowledgments
We thank the referee for their comments and feedback. We thank Andrew Casey and Anna Frebel for co-developing the latest version of SMH. We thank Chris Garling for useful discussions. APJ was supported by NASA through Hubble Fellowship grant HST-HF2-51393.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. W.C. gratefully acknowledges support from a Gruber Science Fellowship at Yale University. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2139841. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. AC is supported by a Brinson Prize Fellowship at UChicago/KICP.
@@
astropy (Astropy Collaboration et al., 2013, 2018), matplotlib (Hunter, 2007), NumPy (Walt et al., 2011), iPython (Pérez & Granger, 2007), SciPy (Virtanen et al., 2020) corner.py (Foreman-Mackey, 2016), emcee (Foreman-Mackey et al., 2013), gala (Price-Whelan, 2017), galpy222222 http://github.com/jobovy/galpy (Bovy, 2015), CarPy (Kelson, 2003), MOOG (Sneden, 1973; Sobeck et al., 2011) SMH (Casey, 2014) pandas (pandas development team, 2020), seaborn (Waskom et al., 2016), pocoMC (Karamanis et al., 2022b, a), NOIRLab IRAF (Tody, 1986, 1993; Fitzpatrick et al., 2024)
@@ Appendix B Literature references for figures
Our references are drawn from the Local Volume Database (Pace, 2024)232323Release 1.0.5 https://github.com/apace7/local_volume_database. References for UFDs including the LMC systems (Bellazzini et al., 2005; Belokurov et al., 2007; Bhardwaj et al., 2024; Boettcher et al., 2013; Bruce et al., 2023; Cantu et al., 2021; Carlin et al., 2009, 2017; Carlin & Sand, 2018; Casey et al., 2025; Cerny et al., 2021a, b, 2023b, 2023a, 2025; Chiti et al., 2021, 2022, 2023; Cicuéndez et al., 2018; Correnti et al., 2009; Crnojević et al., 2016; Dall’Ora et al., 2006, 2012; Drlica-Wagner et al., 2015; Fritz et al., 2019; Garofalo et al., 2013, 2025; Greco et al., 2008; Hansen et al., 2024; Heiger et al., 2024; Homma et al., 2018, 2019, 2024; Jenkins et al., 2021; Ji et al., 2021; Karczmarek et al., 2015; Kim & Jerjen, 2015a; Kim et al., 2016b; Kirby et al., 2013a, 2015, 2017; Koposov et al., 2011, 2015a, 2015b, 2018; Kuehn et al., 2008; Lee et al., 2009; Li et al., 2017, 2018a; Longeard et al., 2018; Martínez-Vázquez et al., 2015; Mart´ınez-Vázquez et al., 2019, 2021a, 2021b; Mateo et al., 2008; McConnachie, 2012; Medina et al., 2018; Moskowitz & Walker, 2020; Muñoz et al., 2018b; Musella et al., 2009; Mutlu-Pakdil et al., 2018, 2020; Oakes et al., 2022; Pace et al., 2020; Richstein et al., 2022, 2024; Sand et al., 2012; Simon & Geha, 2007; Simon et al., 2011, 2015, 2017; Simon, 2019; Simon et al., 2020; Smith et al., 2023; Spencer et al., 2017, 2018; Stetson et al., 2014; Tan et al., 2025; Torrealba et al., 2016a, b, 2018; Vivas et al., 2016, 2020, 2022; Walker et al., 2009, 2015a, 2015b; Walsh et al., 2008; Wang et al., 2019; Willman et al., 2006, 2011). References for ambiguous systems (Balbinot et al., 2013; Cerny et al., 2023a, c; Conn et al., 2018; Gatto et al., 2022; Homma et al., 2019; Kim & Jerjen, 2015b; Kim et al., 2015, 2016a; Luque et al., 2018; Martin et al., 2016; Mau et al., 2019, 2020; Muñoz et al., 2018b; Simon et al., 2024; Smith et al., 2024; Torrealba et al., 2019). References for the GC systems (Alonso-Garc´ıa et al., 2025; Bandyopadhyay et al., 2025; Baumgardt & Hilker, 2018; Baumgardt et al., 2020; Baumgardt & Vasiliev, 2021; Bradford et al., 2011; Çalışkan et al., 2012; Carballo-Bello et al., 2016; Crociati et al., 2023; Da Costa et al., 1992; Deras et al., 2023; Fanelli et al., 2024; Frelijj et al., 2025; Geisler et al., 2023; Gieles et al., 2021; González-D´ıaz et al., 2023; Hamren et al., 2013; Hankey & Cole, 2011; Harris, 1996; Henao et al., 2025; Johnson et al., 2018; Joo et al., 2019; Jordi et al., 2009; Kirby et al., 2015; Kobulnicky et al., 2005; Koch et al., 2009; Kunder & Butler, 2020; Kunder et al., 2021, 2024; Kurtev et al., 2008; Latour et al., 2025; Leanza et al., 2024; Longeard et al., 2021; Loriga et al., 2025; Montecinos et al., 2021; Muñoz et al., 2018a, b; Pallanca et al., 2023; Richstein et al., 2024; Rosenberg et al., 1998; Saviane et al., 2012; Simpson, 2018; Souza et al., 2021; Strader & Kobulnicky, 2008; Vásquez et al., 2018; Voggel et al., 2016; Weisz et al., 2016).
@@
The chemical abundance references of the UFD stars in Figure 9 are as follows: Boötes I (Gilmore et al., 2013; Frebel et al., 2016; Waller et al., 2023), Boötes II (Ji et al., 2016d), Canes Venatici II (François et al., 2016), Carina II and III (Ji et al., 2020b), Cetus II (Webber et al., 2023), Grus I (Ji et al., 2019), Grus II (Hansen et al., 2020), Coma Berenices and Ursa Major II (Frebel et al., 2010), Hercules (Koch et al., 2008; François et al., 2016), Horologium I (Nagasawa et al., 2018), Leo IV (Simon et al., 2010; François et al., 2016), Pisces II (Spite et al., 2018), Reticulum II (Ji et al., 2016a, c) Segue 1 (Frebel et al., 2014), Segue 2 (Roederer & Kirby, 2014), Triangulum II (Venn et al., 2017; Kirby et al., 2017; Ji et al., 2019), Tucana II (Chiti et al., 2018a, 2023), Tucana III (Hansen et al., 2017; Marshall et al., 2019), Tucana V (Hansen et al., 2024), and Ursa Major I (Waller et al., 2023).